OpenStudy (nbarrera):

MEDAL! Use the quadratic formula to solve x^2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A {–12, 5} B {–11.1, –2.9} C {–5.6, –1.4} D {–3.9, 12.1}

OpenStudy (nbarrera):

@wolf1728

OpenStudy (wolf1728):

a=1 b=7 c=8 x = [-b +-sqroot(b^2 - 4*a*c)] / 2*a x = [-7 +-sqroot(49 -32)] / 2 x = [-7 + sqroot (17)] / 2 x = [-7 - sqroot (17)] / 2 sqroot (17) = 4.1231 x = [-7 + 4.1231] / 2 x = -2.8769/2 x = -1.439 x = [-7 - 4.1231] / 2 x = [-11.1231] / 2 x = -5.56155 Looks like it is C

OpenStudy (nbarrera):

Thx